Given f(x) = x^2 -f if x <-1
3x+2 if x >- -1 , find f(-1)

Answer by stanbon(48558)

Given f(x) = x^2 -f if x < -1
3x+2 if x >= -1 ,
find f(-1).
------------
Since -1>=-1, f(-1) = 3(-1)+2 = -1
================
Cheers,
Stan H.